# Notions of entropy for unitary representations

### v2

@@ 9.7 Further remarks
Just like Theorem A, Theorem C has a number of predecessors in the literature, and so do some of the steps in its proof.
Some of the first that lie beyond the discussion in Subsection 6.3 are Lyons’ calculations in [Lyo05, Lyo10]. These concern the problem of asymptotically counting spanning trees in finite connected graphs using their random weak limits. Lyons shows that this ‘tree entropy’ of the random limit graph is given by a Fuglede–Kadison determinant of its Laplacian. His setting does not require a group action, but it yields results for sofic groups as a special case.
More recent examples are continuations of Deninger’s work on determinantal formulas for the entropies of certain algebraically defined measure-preserving systems, but now for sofic groups and sofic entropy. The first of these was [KL11, Theorem 7.1], due to Kerr and Li. This theme was then taken up by Hayes. In [Hay16], he generalized Deninger’s conjecture to equate the sofic entropy of $\boldsymbol{X}_{f}$ with the Fuglede–Kadison determinant of $f\in\mathbf{Z}[\Gamma]$ whenever $\Gamma$ is sofic. His main result actually allows larger finite matrices over $\mathbf{Z}[\Gamma]$ in place of $f$, which approaches a generalization of the Li–Thom theorem to sofic groups, but Hayes also shows that the full generalization of the Li–Thom theorem is false. See also the alternative proof with further refinements in [Hay21b].
Alongside these papers, Hayes has developed other connections between sofic entropy and representation theory. In [Hay18] he proved that an arbitrary measure-preserving $\Gamma$-system can have completely positive sofic entropy only if its Koopman representation is contained in the infinite inflation of the regular representation. In [Hay17], he computed the sofic entropy of a stationary Gaussian process over $\Gamma$ in terms of the real orthogonal representation that defines its first chaos, generalizing one of the results from [HPT08] for single transformations. The main theorem in [Hay17] is particularly worth comparing with how $\varphi_{\mathrm{sing}}$ and $\varphi_{\mathrm{ac}}$ determine whether $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ equals $\infty$- ∞ in Corollary 9.18 above.
@@
If $\mathfrak{A}=C^{\ast}\Gamma$ but $\Gamma$ is not amenable, then AP entropy is not really a straightforward generalization of the left-hand side of (1.1). And the setting of Theorem C has taken us quite far from our original motivation in the form of Szegő’s theorem. AP entropy is not defined as a limit of finite-dimensional determinants, and finer data such as Verblunsky coefficients have no obvious meaning in this generality.
However, some of those other aspects of Szegő’s theorem make a return in a sequel to the present paper [Aus]. In [Aus], we study uniformly random finite-dimensional representations of free groups, and in particular an annealed version of AP entropy that can be defined using these. This is a representation theoretic analog of Bowen’s annealed sofic entropy (formerly called the ‘f-invariant’) from [Bow10c, Bow10a]. Like annealed sofic entropy, annealed AP entropy admits a precise formula. In studying this entropy and its formula, analogs of many finer features of the theory of orthogonal polynomials on $\mathbf{T}$ re-appear.
For example, sequences of ‘generalized Verblunsky coefficients’ can be used to parametrize positive definite functions over free groups. Rather than depending for their definition on a total ordering of the group, these generalized Verblunksy coefficients depend on the fact that free groups have tree-like Cayley graphs. One of the first main theorems about annealed AP entropy is a formula for it as an infinite series in terms of those coefficients. This can be regarded as a much more complete analog of Szegő’s limit theorem for positive definite functions on free groups, with one entirely new feature: an additional term called ‘zeroth-order’ entropy that reflects the non-amenability of the groups.
We leave further details to [Aus], which also develops applications to large deviations for tuples of random matrices.

### v3

@@ 9.7 Further remarks
Just like Theorem A, Theorem C has a number of predecessors in the literature. Some of the first that lie beyond the discussion in Subsection 6.3 are Lyons’ calculations in [110, 111]. These concern the problem of asymptotically counting spanning trees along sequences of finite connected graphs using their random weak limits. Lyons shows that this ‘tree entropy’ of the random limit graph is given by a Fuglede–Kadison determinant of its Laplacian. His setting does not require a group action, but it yields results for sofic groups as a special case.
More recent examples are continuations of Deninger’s work on determinantal formulas for the entropies of certain algebraically defined measure-preserving systems, but now for sofic groups and sofic entropy. The first of these was due to Kerr and Li [95, Theorem 7.1]. This theme was then taken up by Hayes. In [80], he generalized Deninger’s conjecture to equate the sofic entropy of $\boldsymbol{X}_{f}$ with the Fuglede–Kadison determinant of $f\in\mathbf{Z}[\Gamma]$ whenever $\Gamma$ is sofic. His main result actually allows larger finite matrices over $\mathbf{Z}[\Gamma]$ in place of $f$. In this respect it approaches a generalization of the Li–Thom theorem from [104] to sofic groups, but Hayes also shows that the full generalization of the Li–Thom theorem is false. See also the alternative proof with further refinements in [84].
Alongside those papers, Hayes has developed other connections between sofic entropy and representation theory. In [82] he proved that an arbitrary measure-preserving $\Gamma$-system can have completely positive sofic entropy only if its Koopman representation is contained in the infinite inflation of the regular representation. In [81], he computed the sofic entropy of a stationary Gaussian process over $\Gamma$ in terms of the real orthogonal representation that defines its first chaos, generalizing one of the results from [78] for single transformations. The main theorem in [81] is worth comparing with the way in which $\varphi_{\mathrm{sing}}$ and $\varphi_{\mathrm{ac}}$ determine whether $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ equals $\infty$- ∞ in Corollary 9.19 above.
@@
The setting of Theorem C has taken us quite far from our original motivation in the form of Szegő’s theorem. AP entropy is not defined as a limit of finite-dimensional determinants, and related data such as Verblunsky coefficients have no obvious meaning in this generality.
However, some of those finer aspects of Szegő’s theorem make a return in a sequel to the present paper [10]. That paper studies random AP sequences, and in particular an annealed version of AP entropy that can be defined using these. In the special case of uniformly random finite-dimensional representations of free groups, this is a representation theoretic analog of Bowen’s annealed sofic entropy (formerly called the ‘f-invariant’) from [27, 24]. Like annealed sofic entropy, this instance of annealed AP entropy admits a precise formula.
In studying this entropy and its formula, many features of the theory of orthogonal polynomials on $\mathbf{T}$ reappear. For example, sequences of ‘generalized Verblunsky coefficients’ can be used to parametrize positive definite functions over free groups. Rather than requiring a total ordering of the group, these generalized Verblunksy coefficients depend on the fact that free groups have tree-like Cayley graphs. One of the first main theorems about annealed AP entropy is a formula for it as an infinite series in terms of those coefficients. This can be regarded as a much more complete analog of Szegő’s limit theorem for positive definite functions on free groups, with one entirely new feature: an additional term called ‘zeroth-order’ entropy that reflects the non-amenability of the groups.
We leave further details to [10], which also develops applications to large deviations for tuples of random matrices.